OpenStudy (anonymous):

Find the points on the given curve where the tangent line is horizontal or vertical r=e^(theta)

OpenStudy (anonymous):

ur soo pretty

OpenStudy (ybarrap):

$$ r=e^\theta $$ using polar coordinates $$ x=r\cos(\theta)=e^\theta\cos\theta\\ y=r\sin(\theta) =e^\theta\sin\theta $$ Horizontal tangents occur where \(\cfrac{dy}{dx}=0\) and vertical tangents occur where \(\cfrac{dy}{dx} \to \infty \implies dx \to 0\). $$ \cfrac{dy}{dx}=\cfrac{\cfrac{dy}{d\theta}}{\cfrac{dx}{d\theta}}\\ \cfrac{dy}{dx}=\cfrac{e^\theta\cos\theta +e^\theta\sin\theta}{-e^\theta\sin\theta+e^\theta\cos\theta} $$ So horizontal tangents occur where $$ e^\theta\cos\theta +e^\theta\sin\theta=0\\ \implies \theta=\cfrac{-\pi}{4} $$ Vertical tangents occur where $$ -e^\theta\sin\theta+e^\theta\cos\theta=0\\ \implies \theta=\cfrac{\pi}{4} $$ So, there is a horizontal tangent at $$ <-\cfrac{\pi}{4},e^{\left(-\cfrac{\pi}{4}\right )}> $$ and a vertical tangent at $$ <\cfrac{\pi}{4},e^{\left(\cfrac{\pi}{4}\right )}> $$ Which is the answer to your question. Note that in rectangular coordinates, the horizontal an vertical tangents occur at $$ <e^{-\left(\cfrac{\pi}{4}\right )}\cos\cfrac{\pi}{4},e^{-\left(\cfrac{\pi}{4}\right )}\sin\cfrac{\pi}{4}>=<0.32,-0.32> $$ and $$ <e^{\left(\cfrac{\pi}{4}\right )}\cos\cfrac{\pi}{4},e^{\left(\cfrac{\pi}{4}\right )}\sin\cfrac{\pi}{4}>=<1.55,1.55> $$ respectively. Here is a plot of your equation: http://www.wolframalpha.com/input/?i=r%3De^theta If you enable interactivity, you can confirm roughly our work here. Does all this make sense?

OpenStudy (anonymous):

Yes, thank you!